Group Selection Page HELP Controls Group ITDM Vacuum WP78 FXE SPB HED SQS SXP PSPO XO TS Migrated2Zulip Archived2GitlabPages Archived2PDF MigratedByGroup
SCS Photon Commissioning Team XRO DOC WP74 MID XO_mig XO_Operation SQS Migrated2Zulip DET Sample Environment SPB_mig
Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 92     Entry time: 16 Dec 2019, 09:28
Author: Adriano Violante 
Location: General 
Severity: Information 
Subject: Report after dispatch meeting 09.00 

FXE:

  • uTCA update has been completed by EEE on Friday evening. Information will be passed to controls for the finalization.

SCS

  • The interlock system for the big sliding door has to be fixed. JTD is following up the issue with TS and SRP.
  • MT will talk to VAC to clarify how to go on with the train picker integration in the tunnel.

SQS

  • uTCA has been updated.
  • A channel in a digitizer is bisbehaving. The issue is being fixed.
  • Steinmeyer stages will be commissioned today with MS.

MID:

  • SDL tests ongoing
  • Hexapod controllers will be moved to the rackroom. Cables have to be pulled and only after the position of the controllers can be chosen.
  • Loop12 will be updated today by EEE

 EEE:

  • uTCA puppet migration: Instr. done. Tunnels and balconies probably will be updated at the beginning of next year.
  • SASE1 and 3 EPS loop will be updated probably today.
  • PLC in SPB with issues is being worked on.
  • Migration to MC2 motors has been agreed with controls, XRO, XPD. in XTD 6 XPD the upgrade will be done in KW2. For the rest a plan has to be made.

 XPD:

  • Some issue with the font size in the new Karabo GUI. Controls will have a look.

 Controls:

  • Karabo updates ongoing. SASE3 today until 13:00. SA1, 2 on Wed.
  • MC2 motors updates has been started also in FXE. 
  • SASE3: meeting with Lasermet for the integration of interlock signals in Karabo. Controls has handed over the scenes to the instruments that will optimize them.
  • firmware of the power supply for the AGIPD has to be updated urgently (by Wed.). CY sent a mail to AM. If no reply the power supplies will be switched off for the update.
ELOG V3.1.4-7c3fd00